OpenStudy (anonymous):

Use part I of the Fundamental Theorem of Calculus to find the derivative of h(x) = \int_{-5}^{\sin(x)} (\cos(t^4)+t) dt

OpenStudy (tkhunny):

That would be... \(\int_{a(t)}^{b(t)}f(t)\;dt = f(b(t)) - a(b(t))\) THAT "Fundamental Theorem"?

OpenStudy (anonymous):

I believe so but I thought that was the formula for finding the antiderivative and this problem asks for the derivative

OpenStudy (amistre64):

what is the derivative of F(b(t)) ?

OpenStudy (amistre64):

and that was not the formula for finding an antiderivative ....

OpenStudy (anonymous):

sinx*cos(x^4)+1?

OpenStudy (amistre64):

more generally the derivative of F(b(x)) since our limits are functions of x, is an application of the chain rule: F'(b(x)) * b'(x)

OpenStudy (amistre64):

\[\int_{a(x)}^{b(x)}f(t)~dt=F(b(x))-F(a(x))\] \[\frac d{dx}\int_{a(x)}^{b(x)}f(t)~dt=\frac d{dx}F(b(x))-\frac d{dx}F(a(x))\]

OpenStudy (amistre64):

the derivative of F is already stated in the integral ... as f(t)

OpenStudy (amistre64):

does this make sense?

OpenStudy (anonymous):

so the derivative of F is cos(x^4)+x?

OpenStudy (amistre64):

F'(b(x)) = f(b(x)) = cos((b(x))^4) + b(x)

OpenStudy (amistre64):

the chain rule then multiplies that by the derivative of b(x)

OpenStudy (anonymous):

cos*x(cos((sin(x))^4) + sin(x))?

OpenStudy (anonymous):

cosx*(cos((sin(x))^4) + sin(x)) Is What I meant

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

and F(a(x)) = F(-5) and the derivative of a(x) = 0 so that term 'goes away'

OpenStudy (tkhunny):

We PLAY LIKE we can find the anti-derivative and then we use it. This is an awesome result.

OpenStudy (tkhunny):

Whoops. Just noticed type in first response. f(b(t)) not a(b(t)) Of course, amistre64 got it right all along.

OpenStudy (amistre64):

other than an exercise in futility :) i do not see any practical application of taking the derivative of the integral

OpenStudy (anonymous):

So when ever I have a problem like this I will want to first plug in b(x) for the variable and then if b(x) is something more complex then a single variable I will multiply that new equation by the derivative of b(x). Is this correct?

OpenStudy (amistre64):

if that is how you picture the chain rule in action ... yes

OpenStudy (amistre64):

you will want to apply both b(x) and a(x), both limits. it just so happened in this case that a(x) = -5, so the derivative of a=0 F'(a(x)) * a'(x) f(-5) * 0 = 0

OpenStudy (anonymous):

and then subtract F'(b(x))*b'(x) from F'(a(x))*a'(x)?

OpenStudy (amistre64):

other way about \[\frac d{dx}[F(b)-F(a)]=f(b)b'-f(a)a'\]

OpenStudy (anonymous):

Oh. Wow. Okay. I think I am starting to get it. Thank you so much amistre64!

OpenStudy (amistre64):

youre welcome